# Premerger phenomena in neutron-star binary coalescences

### v2

@@ III.1.1 Supporting stress
In elasticity theory, the extent to which a solid can withstand stress is mathematically encapsulated by the so-called Lamé coefficients relating stress and strain [405]. At a linear level, these are just the shear and bulk modulii, the latter of which is expected to be dynamically negligible in the crust [82]. Given that the majority of work regarding oscillations or restorative forces in the crust are discussed at a linear level, and the difficulty of microphysical calculations, little is known about the higher-order coefficients in the crust. We do not discuss nonlinear elasticity further here [though see 27, 360].
Still, the shear modulus, $\mu$, is the leading-order quantity relevant for elastically-supported oscillations and stress support. Even though there really are multiple shear modulii which depend on the shape of nuclei [311], we neglect such complications (e.g. related to the possibility of nuclear “pasta”) and assume a single contributor. For the standard (i.e. spherical-nuclei) shear modulus, the often-quoted expression valid at low temperatures comes from Strohmayer et al. 369, viz.
@@
which is proportional to ${\Gamma}$. More sophisticated variants can also be found, such as that due to Horowitz and Hughto 192 who fit results to molecular dynamics simulations,
@@
Next-to-leading-order temperature corrections are discussed by Baiko 40 and others, as are various physical corrections to these formulae [again see 82]. At the linear (i.e. Hookean) level, $\mu$ is the proportionality factor relating (shear) stress, $\boldsymbol{\sigma}$, to the elastic strain, $\boldsymbol{s}^{\rm el}$, viz. [e.g. 52]
@@
In GR, arriving at a similar relationship is rather more involved, though can achieved through the Carter and Quintana 77 relations defining the elastic stress tensor through a Lie derivative. A modern discussion on relativistic elasticity can be found in Andersson and Comer 27. The shear stress can be related to the Lagrangian eigenfunction associated with generic motions (see Sec. IV.1.1) through
@@
where index symmetry is manifest. The perturbed metric and Christoffel symbols also appear in the GR variant of (6). Although often the Maxwell stress is used directly to define $\boldsymbol{\sigma}$ in works involving magnetar events, this is strictly-speaking invalid: one must instead model the perturbative viscoelastic response to magnetic pressues, which tends to induce significantly lower strains than the Maxwell terms alone [see 215, 69, for recent emphases].
@@ III.2 Breaking strain
Table 1: Estimates for the breaking strain, $\sigma_{\rm max}$, obtained from the literature, in ascending order according to publication date. This list is not exhaustive, though roughly illustrates the plausible range of values depending on assumptions made on crustal microphysics. The parameter $s\approx 0.185Z^{1/3}c/v_{\rm F}$ has been introduced by Kozhberov and Yakovlev 223 for electron Fermi velocity $v_{\rm F}$.
@@
Imperfect (alkali) metals $\gtrsim 10^{-5}$ Smoluchowski 359
Perfect one-component crystal $10^{-2}\lesssim\sigma_{\rm max}\lesssim 10^{-1}$ Kittel 210
Li crystals ($57\%$ Mg, $T\lesssim 100$ K) $10^{-5}\lesssim\sigma_{\rm max}\lesssim 10^{-2}$ Ruderman 342
Energy and event rates of magnetar flares $\lesssim 10^{-3}$ Thompson and Duncan 399
Perfect, defective, and poly-crystal ($\Gamma\sim 834$) $\sim 0.1$ Horowitz and Kadau 193
Perfect body-centred cubic crystal $\left(0.0195-\frac{1.27}{\Gamma-71}\right)n_{i}\frac{Z^{2}e^{2}}{a_{i}}$ Chugunov and Horowitz 90
Pure and imperfect crystals (various compositions) $\gtrsim 0.1$ Hoffman and Heyl 191
Maximum strain set by spin limit ($\lesssim 1$ kHz) $0.008\lesssim\sigma_{\rm max}\lesssim 0.089$ Fattoyev et al. 138
@@
Idealised nuclear pasta ($T=1$ MeV) $\sim 0.3$ Caplan et al. 75
Deformed mono-crystals $\approx 0.02\frac{\sqrt{1+1.451s^{2}}}{1+0.755s^{2}}$ Kozhberov and Yakovlev 223
Multi-ion (strongly ordered) crystal ($T\ll$ MeV) $0.02\lesssim\sigma_{\rm max}\lesssim 0.08$ Kozhberov 222
@@
One anticipates that for ${\Gamma}\gtrsim 175$ the outer layers (except for the very outer layers being oceans and atmosphere) of the neutron star will solidify an elastic solid that can support stress up until a point where a “failure” event occurs. Such stresses can develop through multiple channels, being a general mass quadrupole moment or “mountain” [e.g. 389, 159], deformations due to gradual spindown [48, 206], differential rotation between the crust and interior neutron superfluid [i.e. spherical Couette flow; 341], magnetic field evolution [e.g. 377], or resonant pulsations (see Sec. VI). What this threshold is — the topic of this subsection — has important implications for a variety of phenomena therefore [e.g. 398, 399, 240].
The reason for quotations around the word failure above is that exactly how an overstraining event manifests in the crust is not well understood. As discussed in Sec. VI.3, this may have applications for precursors also. The simplest type of picture one may have of failure is that of a brittle material. The elastic maximum is breached, and suddenly the region “cracks” like glass. In this way one can envision an immediate and large release of magnetoelastic energy: field lines that were once held fixed (cf. Alfvén’s flux-freezing theorem) are now free to reconnect, as the stress holding everything together falters. Jones 202 and other authors have argued against this picture, effectively because the hydrostatic pressure in the crust exceeds the shear modulus by $\gtrsim 2$ orders of magnitude, inhibiting the formation of true voids. It is likely that a more realistic picture is that of plastic deformations: the crust becomes overstrained and undergoes a permanent but not immutable deformation [e.g. 246]. In this case, heat is released as the region fails and twist is injected into the magnetosphere via plastic motions, which prime it for reconnection and magnetic eruptions leading to energy release.
The critical strain, $\sigma_{\rm max}$, has been estimated through a variety of different techniques and approximations over the last $\sim 50$ years, as collated in Table 1. It is a very difficult problem, in general, to estimate global features of the crust via molecular dynamics or other simulations, which are inherently local. For instance, the simulations of Horowitz and Kadau 193 apply for $\sim 10^{11}$ femtometers of material. As emphasised by Kerin and Melatos 206, while the former authors found a global failure mechanism once the critical strain of $\sigma_{\rm max}\sim 0.1$ is reached, it is probable that in a real crust the failure mechanism differs because of lattice dislocations, grain boundaries, permanent or temporary deformities due to previous failures, and other mesoscopic imperfections.
@@
Another mechanism, perhaps more physical as discussed by Chugunov and Horowitz 90 and others, is the Zhurkov model [405]. The main stipulation is that thermodynamic fluctuations should exceed some threshold energy in order for the breaking to occur. It is arguably more realistic since it accounts for the fact that stress is applied over a finite time interval, which leads to a more gradual deformation of the material in question, rather than in an abrupt sense as predicted by (7). Nevertheless, because of its simplicity, the von Mises criterion is often used in the literature [e.g. 240, 381, 377] and is that which we adopt in this Review.
@@ III.2.1 Mountains
Another way in which the critical breaking strain impacts on neutron star predictions is through the maximum mountain size (i.e. mass quadrupole moment or ellipticity). While not especially relevant for premerger phenomena per se, whether or not a star had a history of GW emission could influence evolutionary channels and especially spin (Sec. IV.3). For example, GW radiation-reaction contributes to spindown, and thus it may be that if the maximum mountain size is small [as is predicted by modern approaches; 159], a larger mean spin frequency in late-stage inspirals may apply. This could impact on mode frequency distributions, and the plausibility of late-stage dynamo activity (as described in Sec. VI.5). If a neutron star taking part in a merger already has a mountain through an old pile-up of accreted material [e.g. 267] or some other means, the effective $\sigma_{\rm max}$ necessary to instigate failure would be reduced. Typically though, models assume an initially relaxed (elastic) state for the crust with $\sigma(t=0)=0$.
@@ III.3 Stratification gradients
where $\gamma$ is the adiabatic index associated with the beta equilibrium star, $\gamma=\frac{\epsilon+p}{p}\frac{dp}{d\epsilon}$. The function (parameter) $\tilde{\Gamma}$ — not to be confused with the Coulomb coupling parameter $\Gamma$ — is that associated with the perturbation, generally computed in the “slow-reaction limit” where one assumes that the composition of a perturbed fluid element is frozen, and is related to the sound speed via $c_{s}^{2}=p\tilde{\Gamma}/(\epsilon+p)$. As such, by enforcing that the Lagrangian variation of some particle (proton) fraction is zero, the neighbourhood of perturbed fluid elements are no longer in beta equilibrium and the system supports buoyancy modes [11, 409]. Generally, $\delta$ is both a function of time and space as the star heats and has position-dependent temperature and composition, with the matter such that $\delta\geq 0$ to satisfy the Schwarzschild criterion for convective stability. Compositional impacts are described above, while thermal ones can be estimated following Krüger et al. 224 and others
@@
for particle species $x$, with number density $n_{x}$ and Fermi energy $E^{x}_{F}$, where the sum runs over the species list. Note these latter quantities could also be treated as functions of time if chemical reactions are accounted for. The subscript $T$ indicates a thermal contribution, clearly vanishing as $T\to 0$.
Figure 4: Stratification profile for an isothermal star for the (cold) SLy EOS with a mass of 1.41 $M_{\odot}$, where the crust-core boundary is indicated. A value of $\delta=0.006$ is marked, considered roughly as a canonical average in a number of studies (see text). The curve terminates at the stellar surface, where we do not consider ocean or atmospheric layers. Adapted from Kuan et al. 232.
@@
Within Fig. 5 we normalise the (coordinate) radius by the stellar radius to compare stars with two different EOS. We truncate the domain (i.e. $r/R\geq 0.65$) to show the $\delta$ profile in the outer core and crust only, as the stratification becomes minute at lower radii (cf. Fig. 4). The spiky behaviour near $0.9\lesssim r/R\lesssim 0.95$ corresponds to densities where nuclei crush after reaching their respective saturation values [127]. Although certain types of smoothing are often performed [see, e.g. 175, 318], we did not implement such filtering to the directly-accessible data from CompOSE. The spiky nature of the $\delta$ profile in this case highlights the complications present when trying to calculate realistic $g$-mode spectra and out-of-equilibrium processes when using finite-temperature EOS. Still, the $\delta$ values are quantitatively similar to the case of (9). Strong magnetic fields could also affect the effective stratification [118].
Figure 5: Stratification profile for an isothermal ($T=10^{7}$ K) star with a mass of 1.41 $M_{\odot}$ for the finite temperature SLy4 (blue) and APR (red) EOS accessible within the CompOSE catalogue. The radial coordinate is normalised by the stellar radius for the respective models.
In reality, perturbations of a neutron star will lead to a departure from beta equilibrium, possible changes to the local viscosity, and heat propagation [198, 367, 199, 188]. The local effect of beta gradients can kill off $g$-modes with a frequency lower than the reaction rate as a result of the suppressed compositionally supported buoyancy [33, 98], combined with the fact that heating leads to a shift in mode frequencies (see Sec. IV.5). The influences of diffusive physics in the oscillation spectrum and in the tidal interaction, in the context of coalescing binary neutron stars, are still not well understood. Some recent attempts have been made with simplified models [176, 337, 344]; in particular, Ripley et al. 338 suggested that these effects lead to a measurable phase shift in the GWs for high signal-to-noise-ratio events so that constraints on the bulk viscosity can be placed.
@@ III.4 Superfluidity and superconductivity
Mature neutron stars tend to be relatively cold with $k_{B}T\ll$ MeV. In this context, “cold” means that their thermal energies are well below the corresponding (core) Fermi energies, $10\lesssim E_{F}\lesssim 100$ MeV [noting that the Fermi energy of neutrons is $E_{F}\approx 100\,(\rho/2\rho_{0})^{2/3}$ MeV; Shapiro and Teukolsky 352]. Because of the high degree of degeneracy, the neutrons and protons (and perhaps some exotica) occupying the core of the star are expected to become both superfluid and superconducting within $\lesssim 100$ yr [see 180, for a review]. The exact temperature at which such a transition occurs is both density- and pairing-mechanism dependent and a matter of active research, though is in the neighbourhood of $T_{c}\approx 10^{9}$ K [see Figure 1 in 25, for example]. With the possible exception of some rare dynamical capture events [439, see also Sec. IV.6], stars taking part in a merger should be below this temperature.
Various phases of the crust are also expected to be in such low-resistance states. While simple estimates all bit confirm that that the electrons living in the neutron star crust are not superconducting (the critical temperature is practically zero); by contrast, as the strong interaction has an attractive component and neutrons and protons are fermions, they are expected to form a Cooper-pair condensate at low enough temperatures and thus neutrons in the crust are in fact likely to be superfluid [274, 432].
One important aspect of superfluid+superconducting states in premerger objects concerns $g$-modes, and thus dynamical tides more generally. If the neutrons are superfluid they do not contribute to the buoyancy that other fluid constituents experience following some perturbation, as they are free to “move out of the way” [301]. As described by Kantor and Gusakov 204 and others, the neutron component is thus essentially decoupled from the oscillations and so the mass of some given oscillating fluid element is smaller by a factor that depends on the relative particle abundances. Less inertia for the same force implies a greater oscillation frequency; typically, superfluid $g$-modes have factor $\approx 4$ larger eigenfrequencies than their normal counterparts [442, 25]. This scaling of course depends sensitively on the exact EOS, the presence of temperature gradients, spin, and so on.
@@
Superconductivity has been studied less in the premerger context. This could in principle distort the star significantly away from spherical symmetry [239, which influences the tidal coupling] and shift the mode spectra through magnetic corrections as the Lorentz force scales like $\sim H_{c1}B$ which can be large even if $B$ is relatively small (see Sec. V.4 and Suvorov 375). The lower critical field can be estimated through [see eqution 7 in 161]
@@
for proton fraction, $x_{p}$, and effective (entrained) mass, $m^{\ast}_{p}$. In the cores of some heavy stars under some EOS, (10) could reach $\sim 10^{16}$ G. Nevertheless, in instances where Figures are shown in this work, the impacts of superfluidity and superconductivity have been ignored in calculating mode properties.

### v3

@@ III.1.1 Supporting stress
In elasticity theory, the extent to which a solid can withstand stress is mathematically encapsulated by the so-called Lamé coefficients relating stress and strain [419]. At a linear level, these are just the shear and bulk modulii, the latter of which is expected to be dynamically negligible in the crust [81]. Given that the majority of work regarding oscillations or restorative forces in the crust are discussed at a linear level, and the difficulty of microphysical calculations, little is known about the higher-order coefficients in the crust. We do not discuss nonlinear elasticity further here [though see 27, 374].
Still, the shear modulus, $\mu$, is the leading-order quantity relevant for elastically-supported oscillations and stress support. Even though there really are multiple shear modulii which depend on the shape of nuclei [324], we neglect such complications (e.g. related to the possibility of nuclear “pasta”) and assume a single contributor. For the standard (i.e. spherical-nuclei) shear modulus, the often-quoted expression valid at low temperatures comes from Strohmayer et al. [383], viz.
@@
which is proportional to ${\Gamma}$. More sophisticated variants can also be found, such as that due to Horowitz and Hughto [199] who fit results to molecular dynamics simulations,
@@
Next-to-leading-order temperature corrections are discussed by Baiko [40] and others, as are various physical corrections to these formulae [again see 81]. At the linear (i.e. Hookean) level, $\mu$ is the proportionality factor relating (shear) stress, $\boldsymbol{\sigma}$, to the elastic strain, $\boldsymbol{s}^{\rm el}$, viz. [e.g. 51]
@@
In GR, arriving at a similar relationship is rather more involved, though be can achieved through the Carter and Quintana [76] relations defining the elastic stress tensor through a Lie derivative. A modern discussion on relativistic elasticity can be found in Andersson and Comer [27]. The shear stress can be related to the Lagrangian eigenfunction associated with generic motions (see Sec. IV.1.1) through
@@
where index symmetry is manifest. The perturbed metric and Christoffel symbols also appear in the GR variant of (6). Although often the Maxwell stress is used directly to define $\boldsymbol{\sigma}$ in works involving magnetar events, this is strictly-speaking invalid: one must instead model the perturbative viscoelastic response to magnetic pressues, which tends to induce significantly lower strains than the Maxwell terms alone [see 223, 68, for recent emphases].
@@ III.2 Breaking strain
Table 1: Estimates for the breaking strain, $\sigma_{\rm max}$, obtained from the literature, in ascending order according to publication date. This list is not exhaustive, though roughly illustrates the plausible range of values depending on assumptions made on crustal microphysics. The parameter $s\approx 0.185Z^{1/3}c/v_{\rm F}$ has been introduced by Kozhberov and Yakovlev 231 for electron Fermi velocity $v_{\rm F}$.
@@
Imperfect (alkali) metals $\gtrsim 10^{-5}$ Smoluchowski 373
Perfect one-component crystal $10^{-2}\lesssim\sigma_{\rm max}\lesssim 10^{-1}$ Kittel 217
Li crystals ($57\%$ Mg, $T\lesssim 100$ K) $10^{-5}\lesssim\sigma_{\rm max}\lesssim 10^{-2}$ Ruderman 356
Energy and event rates of magnetar flares $\lesssim 10^{-3}$ Thompson and Duncan 413
Perfect, defective, and poly-crystal ($\Gamma\sim 834$) $\sim 0.1$ Horowitz and Kadau 200
Perfect body-centred cubic crystal $\left(0.0195-\frac{1.27}{\Gamma-71}\right)n_{i}\frac{Z^{2}e^{2}}{a_{i}}$ Chugunov and Horowitz 89
Pure and imperfect crystals (various compositions) $\gtrsim 0.1$ Hoffman and Heyl 198
Maximum strain set by spin limit ($\lesssim 1$ kHz) $0.008\lesssim\sigma_{\rm max}\lesssim 0.089$ Fattoyev et al. 141
@@
Idealised nuclear pasta ($T=1$ MeV) $\sim 0.3$ Caplan et al. 74
Deformed mono-crystals $\approx 0.02\frac{\sqrt{1+1.451s^{2}}}{1+0.755s^{2}}$ Kozhberov and Yakovlev 231
Multi-ion (strongly ordered) crystal ($T\ll$ MeV) $0.02\lesssim\sigma_{\rm max}\lesssim 0.08$ Kozhberov 230
@@
One anticipates that for ${\Gamma}\gtrsim 175$ the outer layers (except for the very outer layers being oceans and atmosphere) of the neutron star will solidify an elastic solid that can support stress up until a point where a “failure” event occurs. Such stresses can develop through multiple channels, being a general mass quadrupole moment or “mountain” [e.g. 403, 164], deformations due to gradual spindown [48, 214], differential rotation between the crust and interior neutron superfluid [i.e. spherical Couette flow; 355], magnetic field evolution [e.g. 391], or resonant pulsations (see Sec. VI). What this threshold is — the topic of this subsection — has important implications for a variety of phenomena therefore [e.g. 412, 413, 249].
The reason for quotations around the word failure above is that exactly how an overstraining event manifests in the crust is not well understood. As discussed in Sec. VI.3, this may have applications for precursors also. The simplest type of picture one may have of failure is that of a brittle material. The elastic maximum is breached, and suddenly the region “cracks” like glass. In this way one can envision an immediate and large release of magnetoelastic energy: field lines that were once held fixed (cf. Alfvén’s flux-freezing theorem) are now free to reconnect, as the stress holding everything together falters. Jones 209 and other authors have argued against this picture, effectively because the hydrostatic pressure in the crust exceeds the shear modulus by $\gtrsim 2$ orders of magnitude, inhibiting the formation of true voids. It is likely that a more realistic picture is that of plastic deformations: the crust becomes overstrained and undergoes a permanent but not immutable deformation [e.g. 255]. In this case, heat is released as the region fails and twist is injected into the magnetosphere via plastic motions, which prime it for reconnection and magnetic eruptions leading to energy release.
The critical strain, $\sigma_{\rm max}$, has been estimated through a variety of different techniques and approximations over the last $\sim 50$ years, as collated in Table 1. It is a very difficult problem, in general, to estimate global features of the crust via molecular dynamics or other simulations, which are inherently local. For instance, the simulations of Horowitz and Kadau 200 apply for $\sim 10^{11}$ femtometers of material. As emphasised by Kerin and Melatos 214, while the former authors found a global failure mechanism once the critical strain of $\sigma_{\rm max}\sim 0.1$ is reached, it is probable that in a real crust the failure mechanism differs because of lattice dislocations, grain boundaries, permanent or temporary deformities due to previous failures, and other mesoscopic imperfections.
@@
Another mechanism, perhaps more physical as discussed by Chugunov and Horowitz 89 and others, is the Zhurkov model [419]. The main stipulation is that thermodynamic fluctuations should exceed some threshold energy in order for the breaking to occur. It is arguably more realistic since it accounts for the fact that stress is applied over a finite time interval, which leads to a more gradual deformation of the material in question, rather than in an abrupt sense as predicted by (7). Nevertheless, because of its simplicity, the von Mises criterion is often used in the literature [e.g. 249, 395, 391] and is that which we adopt in this Review.
@@ III.2.1 Mountains
Another way in which the critical breaking strain impacts on neutron star predictions is through the maximum mountain size (i.e. mass quadrupole moment or ellipticity). While not especially relevant for premerger phenomena per se, whether or not a star had a history of GW emission could influence evolutionary channels and especially spin (Sec. IV.3). For example, GW radiation-reaction contributes to spindown, and thus it may be that if the maximum mountain size is small [as is predicted by modern approaches; 164], a larger mean spin frequency in late-stage inspirals may apply. This could impact on mode frequency distributions, and the plausibility of late-stage dynamo activity (as described in Sec. VI.5). If a neutron star taking part in a merger already has a mountain through an old pile-up of accreted material [e.g. 278] or some other means, the effective $\sigma_{\rm max}$ necessary to instigate failure would be reduced. Typically though, models assume an initially relaxed (elastic) state for the crust with $\sigma(t=0)=0$.
@@ III.3 Stratification gradients
where $\gamma$ is the adiabatic index associated with the beta equilibrium star, $\gamma=\frac{\epsilon+p}{p}\frac{dp}{d\epsilon}$. The function (parameter) $\tilde{\Gamma}$ — not to be confused with the Coulomb coupling parameter $\Gamma$ — is that associated with the perturbation, generally computed in the “slow-reaction limit” where one assumes that the composition of a perturbed fluid element is frozen, and is related to the sound speed via $c_{s}^{2}=p\tilde{\Gamma}/(\epsilon+p)$. As such, by enforcing that the Lagrangian variation of some particle (proton) fraction is zero, the neighbourhood of perturbed fluid elements are no longer in beta equilibrium and the system supports buoyancy modes [11, 423]. Generally, $\delta$ is both a function of time and space as the star heats and has position-dependent temperature and composition, with the matter such that $\delta\geq 0$ to satisfy the Schwarzschild criterion for convective stability. Compositional impacts are described above, while thermal ones can be estimated following Krüger et al. 233 and others
@@
for particle species $x$, with number density $n_{x}$ and Fermi energy $E^{x}_{F}$, where the sum runs over the species list. Note these latter quantities could also be treated as functions of time if chemical reactions are accounted for. The subscript $T$ indicates the thermal contribution, which clearly vanishes as $T\to 0$.
Figure 4: Stratification profile for an isothermal star for the (cold) SLy EOS with a mass of 1.41 $M_{\odot}$, where the crust-core boundary is indicated. A value of $\delta=0.006$ is marked, considered roughly as a canonical average in a number of studies (see text). The curve terminates at the stellar surface, where we do not consider ocean or atmospheric layers. Adapted from Kuan et al. 241.
@@
Within Fig. 5 we normalise the (coordinate) radius by the stellar radius to compare stars with two different EOS. We truncate the domain (i.e. $r/R\geq 0.65$) to show the $\delta$ profile in the outer core and crust only, as the stratification becomes minute at lower radii (cf. Fig. 4). The spiky behaviour near $0.9\lesssim r/R\lesssim 0.95$ corresponds to densities where nuclei crush after reaching their respective saturation values [128]. Although certain types of smoothing are often performed [see, e.g. 181, 331], we did not implement such filtering to the directly-accessible data from CompOSE. A comparison between these two figures effectively demonstrate the validity of the thermal-pressure approximation (9) as compared to the use of a precise, finite-temperature EOS. Strong magnetic fields could also affect the effective stratification [118].
Figure 5: Stratification profile for an isothermal ($T=10^{7}$ K) star with a mass of 1.41 $M_{\odot}$ for the finite temperature SLy4 (blue) and APR (red) EOS accessible within the CompOSE catalogue. The radial coordinate is normalised by the stellar radius for the respective stellar models.
In reality, perturbations of a neutron star will lead to a departure from beta equilibrium, possible changes to the local viscosity, and heat propagation [205, 381, 206, 194]. The local effect of beta gradients can kill off $g$-modes with a frequency lower than the reaction rate as a result of the suppressed compositionally supported buoyancy [33, 97], combined with the fact that heating leads to a shift in mode frequencies (see Sec. IV.5). The influences of diffusive physics in the oscillation spectrum and in the tidal interaction, in the context of coalescing binary neutron stars, are still not well understood. Some recent attempts have been made with simplified models [182, 351, 358]; in particular, Ripley et al. 352 suggested that these effects lead to a measurable phase shift in the GWs for high signal-to-noise-ratio events so that constraints on the bulk viscosity can be placed.
@@ III.4 Superfluidity and superconductivity
Mature neutron stars tend to be relatively cold with $k_{B}T\ll$ MeV. In this context, “cold” means that their thermal energies are well below the corresponding (core) Fermi energies, $10\lesssim E_{F}\lesssim 100$ MeV [noting that the Fermi energy of neutrons is $E_{F}\approx 100\,(\rho/2\rho_{0})^{2/3}$ MeV; [366]]. Because of the high degree of degeneracy, the neutrons and protons (and perhaps some exotica) occupying the core of the star are expected to become both superfluid and superconducting within $\lesssim 100$ yr (see [186] for a review). This was speculated on first by Migdal [285], with the first proper calculations carried out by Ginzburg & Kirzhniz [162] providing support. Since then, more elaborate calculations have been performed; see, for instance, Gandolfi et al. 158, Drissi & Rios 130, Krotscheck, Papakonstantinou, & Wang 232 for recent contributions based on different many-body methods.
The exact temperature at which such a transition occurs is both density- and pairing-mechanism dependent and a matter of active research, though is in the neighbourhood of $T_{c}\approx 10^{9}$ K (see Figure 1 in [25] for example). With the possible exception of some rare dynamical capture events [453, see also Sec. IV.6], stars taking part in a merger should be below this temperature. Various phases of the crust are also expected to be in such low-resistance states. As the strong interaction has an attractive component and neutrons and protons are fermions, they are expected to form a Cooper-pair condensate at low enough temperatures and thus neutrons in the crust are in fact likely to be superfluid [446]; by contrast, simple estimates all but confirm that that the electrons living in the neutron star crust are not superconducting (the critical temperature is practically zero).
One important aspect of superfluid+superconducting states in premerger objects concerns $g$-modes, and thus dynamical tides more generally. If the neutrons are superfluid they do not contribute to the buoyancy that other fluid constituents experience following some perturbation, as they are free to “move out of the way” [314]. As described by Kantor and Gusakov [212] and others, the neutron component is thus essentially decoupled from the oscillations and so the mass of some given oscillating fluid element is smaller by a factor that depends on the relative particle abundances. Less inertia for the same force implies a greater oscillation frequency; typically, superfluid $g$-modes have factor $\approx 4$ larger eigenfrequencies than their normal counterparts [456, 25]. This scaling of course depends sensitively on the exact EOS, the presence of temperature gradients, spin, and so on.
@@
Superconductivity has been studied less in the premerger context. This could in principle distort the star significantly away from spherical symmetry [248] (which influences the tidal coupling) and shift the mode spectra through magnetic corrections as the Lorentz force scales like $\sim H_{c1}B$ which can be large even if $B$ is relatively small (see Sec. V.4 and [389]). The lower critical field can be estimated through [see equation (7) in [166]]
@@
for proton fraction, $x_{p}$, and effective (entrained) mass, $m^{\ast}_{p}$. In the cores of some heavy stars under some EOS, equation (10) could reach $\sim 10^{16}$ G. Nevertheless, in instances where Figures are shown in this work, the impacts of superfluidity and superconductivity have been ignored in calculating mode properties.